You are given three investment alternatives to analyze. The cash flows from these three investments are as follows:

Investment

End of Year     A     B     C

1 $ 1,000  $ 1,000 $ 4,000  

2    2,000   1,000 4,000  

3    3,000   1,000 (4,000)

4    (4,000) 1,000 (4,000)

5    4,000   3,000    14,000  

What is the present value of each of these three investments if the appropriate discount rate is 11 percent? (Round to the nearest cent.)
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

A:

Present value=1000/1.11+2000/1.11^2+3000/1.11^3-4000/1.11^4+4000/1.11^5

=$4456.60(Approx).

B:

Present value=1000/1.11+1000/1.11^2+1000/1.11^3+1000/1.11^4+3000/1.11^5

=$4882.80(Approx).

C:

Present value=4000/1.11+4000/1.11^2-4000/1.11^3-4000/1.11^4+14000/1.11^5

=$9598.72(Approx).